OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (jdoe0001):

test them let us test, say the 1st one 5 , 3 x , y 3x-2y=9 3(5) - 2(3) = 9 15 - 6 =9 9 = 9 \(\large \checkmark\) 3x + 2y = 14 3(5) + 2(3) = 14 15 + 6 = 14 21 \(\ne\) 14 <--- no dice, thus is not that one then check the others

OpenStudy (anonymous):

what the hell that didn't even look close to what I do

OpenStudy (anonymous):

I need help with anther one after this if your up for it just these two

OpenStudy (jdoe0001):

well. that how I'd do it there are other ways, that' d be the simpler one, is all

OpenStudy (jdoe0001):

just post anew, more eyes :)

OpenStudy (anonymous):

is it D i just did the math